sally travels 50m west 120m south find the distance and displacement?
1
Expert's answer
2020-01-13T11:02:49-0500

The distance


"d=d_1+d_2=50+120=170\:\rm m"

The displacement


"{\bf s}={\bf d}_1+{\bf d}_2"

where


"{\bf d}_1=-50{\bf i}+0{\bf j},\; {\bf d}_2=0{\bf i}-120{\bf j}"

Hence


"{\bf s}=-50{\bf i}+0{\bf j}+0{\bf i}-120{\bf j}=-50{\bf i}-120{\bf j}"

The magnitude of a displacement


"s=\sqrt{(-50)^2+(-120)^2}=130\:\rm m"